# Confabulation dynamics in a reservoir computer: Filling in the gaps with untrained attractors

### v2

@@ III.2.2 Reconstruction of $\pazocal{L}$ via period-doubling route to chaos
Above all, Fig. 9 demonstrates the benefit of performing this type of analysis of tracking the changes in these attractors. It provides a strong visual aid to show that when the RC fails to provide any reconstruction of $\pazocal{L}$, there is already an UA waiting for the RC’s state to approach. Furthermore, this is not the only time we see period-doubling bifurcations playing a role in how the RC reconstructs $\pazocal{L}$. In Appendix C we illustrate that period-doubling bifurcations of another Lorenz-like period-1 limit cycle and also a saddle-focus homoclinic orbit result in the reconstruction of $\pazocal{L}$. We would also like to mention that we find no sequence of events where $\hat{\pazocal{L}}$ comes into existence through the well-known sequence of events that unfold when varying the constant term in the $\dot{x}_{2}$ equation in Eq. (26)[45].
@@ IV Task (ii) results
It is clear from Fig. 11 that there is an improvement in the RC’s ability to reconstruct both limit cycles at their respective $b$ values when the trained limit cycles are further apart in $\bm{b}$-space. Panels (a) and (c) show period-4 limit cycles, while panel (e) reveals a similar appearing limit cycle that is actually period-2. Nonetheless, Fig. 12(a)–(b) shows there is a limit to this: if the limit cycles are too far apart, like for $b=\pm 0.5$, then there is a decline in the RC’s ability to reconstruct both limit cycles at their respective $b$ values. The RC fails to reconstruct the period-4 limit cycle and instead produces a chaotic attractor. Further, panels (c)–(d) show what can happen when the limit cycles are too close together in parameter space. For $b=\pm 0.1$ neither limit cycles are reconstructed properly and the RC instead produces period-2 limit cycles at both values of $b$.
@@
The most prominent feature in Fig. 13 (b) is that the continuous transition between the trained period-1 and period-4 limit cycles is not consistent with the period-doubling bifurcation that occurs in Eq. (32). Instead of undergoing a period-doubling bifurcation from the reconstructed period-1 and period-4 limit cycles to a GA in the form of a period-2 limit cycle, the RC generates infinite sequences of period-doubling and halving bifurcation to and from generated chaotic attractors. Fig. 14 (b) illustrates the period-2 limit cycle generated by the RC for $b=0$, which occupies a much larger region of the state space than the corresponding dynamics from Eq. (32).
@@ IV.3 Does adding more data help?
In Fig. 16 we examine how the closed-loop parameter-aware RC in Eq. (20) generates the dynamics between and beyond the three reconstructed limit cycles shown in Fig. 15. More specifically, Fig. 16 show the result of tracking the changes in the dynamics of all attractors generated by Eq. (20) for $b\in\left[0.42,-0.42\right]$. While Fig. 15 shows that the RC achieves a good reconstruction of each limit cycle at their respective $b$ values, the same cannot be said for how the RC generates the dynamics between these limit cycles. While the transition from the period-1 to period-2 limit cycle is relatively consistent with the dynamics of Eq. (32), there is no continuous transition between the period-2 and period-4 limit cycles. The period-2 and period-4 limit cycles both become chaotic when tracking the changes in the dynamics of these attractors for increasing $b$ from 00 and decreasing $b$ from $0.2$, respectively. The chaotic attractor that is born from the reconstructed attractors reaches a point where it can no longer be tracked and thus becomes unstable for $b\in\left(0.12,0.06\right)$. Like before, UA1 fills this gap, while UA1 and UA2 fill in the gap for $b>0.235$. A different scenario from Fig. 13 (c) emerges in Fig. 16 for $b<-0.2$. In addition to UA3, there are two further UAs required to fill in the gaps: the limit cycle UA5 and the mostly chaotic dynamics of UA6.
In Fig. 18 we apply the same procedure to examine how Eq. (20) generates the dynamics between and beyond the five reconstructed limit cycles shown in Fig. 17. Similarly, while Fig. 17 shows that the RC achieves a good reconstruction of each limit cycle at their respective $b$ values, there is a slightly better agreement between the attractors that the RC generates and the dynamics of Eq. (32) than in Fig. 15, aside for the branch of GAs between $b=0.2$ and $0.1$ which shows the RC’s dynamics become chaotic nearby $b=0.18$. Furthermore, there is still no continuous transition between all reconstructed attractors. The periodic and chaotic dynamics of UA7 fills the gap in this case, coexisting with the corresponding GAs for a small range of $b$ values. Just as in many previous cases, UA1 and UA3 fill in the gaps for $b>0.25$ and $b<-0.27$.
@@
Figure 18: Resultant bifurcation diagram from training parameter-aware RC to reconstruct period-1 limit cycle at $b=-0.2$ and $-0.1$, period-2 limit cycle at $b=0.0$ and $0.1$, and period-4 limit cycle at $b=0.2$. Legend indicates the attractors that are tracked.
@@ V Task (iii) results
We first show in Fig. 19 that the closed-loop parameter-aware RC in Eq. (20) achieves a reasonably good reconstruction of the Lorenz ($\pazocal{L}$) and Halvorsen ($\pazocal{H}$) chaotic attractors for $b=\pm 0.2$, ±$\pm 0.3$, and ±$\pm 0.4$, and relatively poor reconstruction for $b=\pm 0.1$, in particular for $\pazocal{L}$. Similar to Fig. 11, Fig. 19 shows that as the attractors are moved further apart in $b$-space then reconstruction of both attractors improves. However, there is also a limit to this as well, comparing Fig. 19 (b) to (c) we see that reconstruction of $\pazocal{H}$ begins to worsen as $b$ is increased from ±$\pm 0.3$ to ±$\pm 0.4$.
@@
Figure 20: Resultant bifurcation diagram from training parameter-aware RC to perform task (iii) with $b=\pm 0.4$, ±$\pm 0.3$, ±$\pm 0.2$, and ±$\pm 0.1$. Legend indicates the attractors that are tracked.
@@
When tracking the changes in $\hat{\pazocal{L}}^{*}$ for decreasing $b$ from $b=0.1$, we find that $\hat{\pazocal{L}}^{*}$ undergoes a series of bifurcations to become a period-1 limit cycle near $b=0.06$ and undergoes a further bifurcation to a fixed point near $b=0.01$. When tracking the changes in $\hat{\pazocal{H}}^{*}$ for increasing $b$ from $b=-0.1$, $\hat{\pazocal{H}}^{*}$ undergoes what appears to be a more abrupt transition from chaotic behaviour to the same fixed point that is found when tracking $\hat{\pazocal{L}}^{*}$. We study these transitions in greater details in the next subsection.
@@
V.3 Examining the continuous transition between $\hat{\pazocal{L}}$ and $\hat{\pazocal{H}}$
To more clearly illustrate that the transition in $\hat{\pazocal{H}}^{*}$ from chaotic to the fixed point dynamics near $b=-0.05$ happens via a bifurcation, in Fig. 21 we re-present Fig. 20 (d) from the perspective of the changes in the local minima of the $x_{3}$-variable. Also, to help further illustrate this bifurcation and the bifurcation from the limit cycle near $b=0.055$, in Figs. 22 and 23, we present snapshots of the RC’s dynamics at different $b$ values before and after these bifurcations take place.
Figure 22: Illustrating transition from $\hat{\pazocal{H}}^{*}$ to fixed point for decreasing $b$ as shown in Figs. 19 (d) and  21.
When looking at Figs. 21 and 22 together we see that there is no direct transition from chaotic attractor to fixed point, but rather the chaotic attractor that exists for $b=-0.056$ transitions to a limit cycle for $b=-0.055$. Prior to this transition, we see in Fig. 22 (c) that significant portions of the trajectory on this chaotic attractor are concentrated around where the limit cycle subsequently emerges. What is common among the limit cycle and chaotic attractor is that the trajectory around each attractor appears to contain a point nearby $(x_{1},x_{3})=(-6,26)$ and the state of the system spends a relatively large duration of time in the vicinity of this point. This suggests these attractors are of the saddle-focus homoclinic orbit type. Furthermore, it is this point (the saddle-focus) which becomes the stable fixed point for $b=-0.054$ shown in Fig. 22 (a).
In Fig. 23 we take a closer look at the transition in $\hat{\pazocal{L}}^{*}$ from Lorenz-like dynamics to the limit cycle near $b=0.06$ in Fig. 21. In Fig. 23 (a) we see that the state of the RC spends much larger durations of time in the left-hand wing of the attractor. The sharp corner in the trajectory nearby $(x_{1},x_{3})=(-1,12)$ indicates there is a saddle nearby, as flow converges towards this point along the $x_{3}$ direction and diverges away from this point along the $x_{1}$ direction. When decreasing $b$ from $0.062$ in Fig. 23 (a) to $0.06$ in Fig. 23 (b), it appears that the Lorenz-like attractor and saddle collide, which results in the right-hand wing disappearing and subsequently results in the period-4 limit cycle seen in Fig. 23 (b). We choose to show the change in the dynamics of this limit cycle from $b=0.06$ in Fig. 23 (b) to $b=0.04$ in Fig. 23 (b) in order to compliment the dynamics shown in Fig. 21, that the limit cycle continues to shrink until it becomes point-like and subsequently becomes the fixed point shown explicitly in Fig. 22 (a), which suggests that a Hopf bifurcation is responsible for this transition.
Figure 23: Illustrating transition from $\hat{\pazocal{L}}^{*}$ to limit cycle for decreasing $b$ as shown in Figs. 19 (d) and  21.
While we do not see regions of bistability between the GAs from the reconstructions of the different limit cycles in task (ii), the results presented in Figs. 20 (d) and 21 indicates that such a scenario is also possible. This leads us to make the following, more general, statement. If the RC fails to interpolate between reconstructed attractors via a sequence of continuous transitions then two options are possible: there is either a region of bistability between GAs that arise from the reconstructed attractors, or there is an UA that coexists with GAs or reconstructed attractors at different parameter settings.
@@ VI.1 Complete confabulation conjecture
Above all, our results suggest that complete confabulations in the form of UAs are not a random occurrence but a by-product of the system being bounded, a system whose state cannot approach infinity. In the RC we study, this boundedness property is induced by the $\tanh{\left(\cdot\right)}$ activation function as it confines the elements of the RC’s state to be valued between $-1$ and $1$. Thus, the state space of the RC is bounded. As a result, when reconstruction fails, the system does not become unstable with its state tending towards infinity, instead there is an UA which is waiting for the system’s state to approach it. In this sense, UAs act as analogues to attractors at infinity, a safety net which prevents the state of the system from approaching infinity and thus may be a desirable feature in certain learning systems that are not bounded. However, since many learning systems share this boundedness property due to the choice of activation function, confabulations similar to UAs may also be viewed as an intrinsic property of these systems.
Extending this logic to biological learning systems, we note that the brain is also subject to hard physiological limits: neurons can only respond to stimuli within a finite dynamical range. This suggests that complete confabulations may appear as a consequence of this biological boundedness. When a bounded learning system fails to encode a memory or generalise between memories, it must ‘fill in’ the gap, resulting in a false, yet sometimes plausible, memory.
In response to our findings and the statements made above, we arrive at the following more general statement on the nature of complete confabulation:
@@
This conjecture offers a unifying principle for understanding the main reason why complete confabulations appear in artificial and biological learning systems. In our future research we will test the validity of the complete confabulation conjecture across different types of RCs, other machine learning paradigms, different tasks, and levels of biological realism.
@@ Appendix A RC design and training parameters
In this paper, M is constructed with an Erdös-Renyi topology where each of the non-zero elements are then replaced with a random number between $-1$ and $1$, the matrix is subsequently scaled to a specific spectral radius, $\rho$. To elaborate, M is designed such that each element is chosen independently to be nonzero with probability $P$ (i.e. sparsity $=P$ or degree $=N/P$) and these nonzero elements are chosen uniformly from $\left(-1,1\right)$. This random sparse matrix is rescaled such that the spectral radius, which is the maximum of the absolute values of its eigenvalues, is $\rho$. This topology is chosen in order to provide the RC with enough dynamical flexibility to solicit multistable dynamics in tasks requiring multifunctionality. Similarly, the input matrix, $\textbf{W}_{in}$, is designed such that each row has only one nonzero randomly assigned element, chosen uniformly from $\left(-1,1\right)$.
@@ 3. C3:
Figure 24: Illustration of decision process in distinguishing between $\hat{\pazocal{L}}$ in (a) and $\tilde{\pazocal{L}}$ in (b). Blue points indicate local maxima that satisfy classification criteria, red point indicate local maxima that do not. Solid black lines indicate a good fit between $\pazocal{L}$ and its reconstruction, dashed black lines indicate a poor fit.
Appendix C Additional reconstruction routes for $\pazocal{L}$

### v3

@@ III.2.1 Reconstruction of $\pazocal{L}$ via period-doubling route to chaos
In Fig. 9, we show that by increasing the parameter $\rho$ of $\textbf{M}_{28}$ from 0, the first time the RC reconstructs a Lorenz-like attractor, it reconstructs $\pazocal{L}$ as a limit cycle. In the inset panel we see that this limit cycle undergoes a period-doubling route to chaos before becoming unstable. To provide further insight, in Fig. 10 we illustrate the dynamics of this limit cycle as it undergoes these bifurcations at different $\rho$ values. As indicated by the arrow in Fig. 10 (e), after this reconstruction becomes unstable the state of the RC then approaches UA1 as this is the only stable state in $\mathbb{P}$ for these $\rho$ values. Prior to becoming unstable, this reconstruction grows in its resemblance to $\pazocal{L}$, which is why our algorithm classifies this as a poor reconstruction of $\pazocal{L}$ rather than an UA. At larger $\rho$ values, we see this poor reconstruction of $\pazocal{L}$ reappear and disappear within different ranges of $\rho$ values for $\rho<0.1$ in a similar way to the events shown in Fig. 21 in Appendix C. Within this range of $\rho$ values we also find another UA, a fixed point that we denote by UA5, which the state of the RC approaches when reconstruction fails for $0.025\lesssim\rho\lesssim 0.03$. For $0.1<\rho<0.325$ we find that the reconstruction exists as a transient and, depending on the value of $\rho$, this transient activity can last for significantly long durations of time. In order to determine whether the RC provides a stable reconstruction of $\pazocal{L}$ within this large range of $\rho$ values, we had to increase $t_{predict}$ to 30,00030,00030 , 000. The RC achieves a good reconstruction of $\pazocal{L}$ for $\rho=0.325$ and maintains this for much larger $\rho$ values.
Above all, Fig. 9 demonstrates the benefit of performing this type of analysis of tracking the changes in these attractors. It provides a strong visual aid to show that when the RC fails to provide any reconstruction of $\pazocal{L}$, there is already an UA waiting for the RC’s state to approach. Furthermore, this is not the only time we see period-doubling bifurcations playing a role in how the RC reconstructs $\pazocal{L}$. In Appendix D we illustrate that period-doubling bifurcations of another Lorenz-like period-1 limit cycle and also a saddle-focus homoclinic orbit result in the reconstruction of $\pazocal{L}$. We would also like to mention that we find no sequence of events where $\hat{\pazocal{L}}$ comes into existence through the well-known sequence of events that unfold when varying the constant term in the $\dot{x}_{2}$ equation in Eq. (26)[48].
@@ IV Task (ii) results
It is clear from Fig. 11 that there is an improvement in the RC’s ability to reconstruct both limit cycles at their respective $b$ values when the trained limit cycles are further apart in $\bm{b}$-space. Panels (a) and (c) show period-4 limit cycles, while panel (e) reveals a similar appearing limit cycle that is actually period-2. Nonetheless, Figs. 12 (a)–(b) shows there is a limit to this: if the limit cycles are too far apart, like for $b=\pm 0.5$, then there is a decline in the RC’s ability to reconstruct both limit cycles at their respective $b$ values. The RC fails to reconstruct the period-4 limit cycle and instead produces a chaotic attractor. Further, panels (c)–(d) show what can happen when the limit cycles are too close together in parameter space. For $b=\pm 0.1$ neither limit cycles are reconstructed properly and the RC instead produces period-2 limit cycles at both values of $b$.
@@
The most prominent feature in Fig. 13 (b) is that the continuous transition between the trained period-1 and period-4 limit cycles is not consistent with the period-doubling bifurcation that occurs in Eq. (32). Instead of undergoing a period-doubling bifurcation from the reconstructed period-1 and period-4 limit cycles to a GA in the form of a period-2 limit cycle, the RC generates infinite sequences of period-doubling and halving bifurcations to and from generated chaotic attractors. Fig. 14 (b) illustrates the period-2 limit cycle generated by the RC for $b=0$, which occupies a much larger region of the state space than the corresponding dynamics from Eq. (32).
@@ IV.3 Does adding more data help?
In Fig. 16 we examine how the closed-loop parameter-aware RC in Eq. (20) generates the dynamics between and beyond the three reconstructed limit cycles shown in Fig. 15. More specifically, Fig. 16 shows the result of tracking the changes in the dynamics of all attractors generated by Eq. (20) for $b\in\left[0.42,-0.42\right]$. While Fig. 15 shows that the RC achieves a good reconstruction of each limit cycle at their respective $b$ values, the same cannot be said for how the RC generates the dynamics between these limit cycles. While the transition from the period-1 to period-2 limit cycle is relatively consistent with the dynamics of Eq. (32), there is no continuous transition between the period-2 and period-4 limit cycles. The period-2 and period-4 limit cycles both become chaotic when tracking the changes in the dynamics of these attractors for increasing $b$ from 0 and decreasing $b$ from 0.20.20.2, respectively. The chaotic attractor that is born from the reconstructed attractors reaches a point where it can no longer be tracked and thus becomes unstable for $b\in\left(0.12,0.06\right)$. Like before, UA1 fills this gap, while UA1 and UA2 fill in the gap for $b>0.235$. A different scenario from Fig. 13 (c) emerges in Fig. 16 for $b<-0.2$. In addition to UA3, there are two further UAs required to fill in the gaps: the limit cycle UA5 and the mostly chaotic dynamics of UA6.
In Fig. 18 we apply the same procedure to examine how Eq. (20) generates the dynamics between and beyond the five reconstructed limit cycles shown in Fig. 17. Similarly, while Fig. 17 shows that the RC achieves a good reconstruction of each limit cycle at their respective $b$ values, there is a slightly better agreement between the attractors that the RC generates and the dynamics of Eq. (32) than in Fig. 15, aside for the branch of GAs between $b=0.2$ and 0.10.10.1 which shows the RC’s dynamics become chaotic nearby $b=0.18$. Furthermore, there is still no continuous transition between all reconstructed attractors. The periodic and chaotic dynamics of UA7 fills the gap in this case, coexisting with the corresponding GAs for a small range of $b$ values. Just as in many previous cases, UA1 and UA3 fill in the gaps for $b>0.25$ and $b<-0.27$.
@@
Figure 18: Resultant bifurcation diagram from training parameter-aware RC to reconstruct period-1 limit cycle at $b=-0.2$ and 0.1-0.1- 0.1, period-2 limit cycle at $b=0.0$ and 0.10.10.1, and period-4 limit cycle at $b=0.2$. Legend indicates the attractors that are tracked.
@@ V Task (iii) results
We first show in Fig. 19 that the closed-loop parameter-aware RC in Eq. (20) achieves a reasonably good reconstruction of the Lorenz ($\pazocal{L}$) and Halvorsen ($\pazocal{H}$) chaotic attractors for $b=\pm 0.2$, ±0.3$\pm$ 0.3± 0.3, and ±0.4$\pm$ 0.4± 0.4, and relatively poor reconstruction for $b=\pm 0.1$, in particular for $\pazocal{L}$. Similar to Fig. 11, Fig. 19 shows that as the attractors are moved further apart in $b$-space then reconstruction of both attractors improves. However, there is also a limit to this as well, comparing Fig. 19 (b) to (c) we see that reconstruction of $\pazocal{H}$ begins to worsen as $b$ is increased from ±0.3$\pm$ 0.3± 0.3 to ±0.4$\pm$ 0.4± 0.4.
@@
Figure 20: Resultant bifurcation diagram from training parameter-aware RC to perform task (iii) with $b=\pm 0.4$, ±0.3$\pm$ 0.3± 0.3, ±0.2$\pm$ 0.2± 0.2, and ±0.1$\pm$ 0.1± 0.1. Legend indicates the attractors that are tracked.
@@
When tracking the changes in $\hat{\pazocal{L}}^{*}$ for decreasing $b$ from $b=0.1$, we find that $\hat{\pazocal{L}}^{*}$ undergoes a series of bifurcations to become a period-1 limit cycle near $b=0.06$ and undergoes a further bifurcation to a fixed point near $b=0.01$. When tracking the changes in $\hat{\pazocal{H}}^{*}$ for increasing $b$ from $b=-0.1$, $\hat{\pazocal{H}}^{*}$ undergoes what appears to be a more abrupt transition from chaotic behaviour to the same fixed point that is found when tracking $\hat{\pazocal{L}}^{*}$. We study these transitions in greater detail in Appendix E.
@@ VI Conclusion
While the results presented in this paper are based on ‘toy examples’ (synthetic data, not real-world data), in the future we intend to conduct follow-up research to the work presented here using real-world data. However, we would like to take the opportunity to highlight investigations of confabulation in learning systems that include a real-world component. For instance, in terms of confabulation in physical RCs, in Terajima et al. [42] the authors conduct an in-depth study of UAs that appear in a multifunctional physical RC in the form of a soft tensegrity robot. More specifically, the authors trained the robot to reconstruct attractors that correspond to two different motions (rapid crawl and shake). However, after further investigation of the robots state space, the authors found that the robot has five additional UAs present in its state space which correspond to completely different motions that it was not trained to exhibit. In terms of confabulation in large language models (LLMs), in Castilho et al. [54] the authors present an interesting study of how different versions of ChatGPT confabulate when used to translate sentences from the English to Irish language (Gaeilge). Interestingly, the authors identify that these LLMs confabulate in a variety of ways by replacing unknown words with made-up words that, for instance, do or do not follow plausible translation rules, or disregard conventions of the Irish alphabet (which does not include certain letters from the English alphabet like ‘v’). In terms of confabulation in the human brain, in El Haj and Larøi [55] the authors present a study of complete confabulations in patients with Alzheimer’s disease, concluding that complete confabulations are correlated with patients that have spatial and temporal disorientation and that momentary confabulations occur more often than complete confabulations which, interestingly and perhaps coincidentally, corresponds with our findings in Fig. 5.
@@ VI.1 Complete confabulation conjecture
Our results suggest that complete confabulations in the form of UAs are not a random occurrence but a by-product of the system being bounded (i.e. a system whose state cannot approach infinity). In the RC we study, this boundedness property is induced by the $\tanh{\left(\cdot\right)}$ activation function as it confines the elements of the RC’s state to be valued between 1-1- 1 and 111. Thus, with both the decay term and bounded activation function the state of the system $\bm{r}(t)$ cannot grow to infinity and the state space of the RC is bounded. As a result, when reconstruction fails, the system does not become unstable with its state tending towards infinity. Instead, since the system cannot approach infinity, it must approach an alterative attractor, which is the UA. In contrast, when reconstruction fails using a so-called ‘next-generation RC’, the system may tend to infinity because there is no guarantee of boundedness in the design of a next-generation RC. Examples of this occurrence can be found in Fig. 1 of Flynn et al. [56].
Here we suggest that the boundedness property of a system may serve as a unifying principle for the intrinsic existence of complete confabulations in artificial and biological learning systems:
@@
Many learning systems share this property of boundedness. The state of most deep neural networks will be bounded due to the use of common activation functions that are either bounded or asymptotically linear [57], such that any bounded input will generate a bounded output. Similarly, biological neurons in the brain are subject to physiological constraints and can only respond to stimuli within a finite dynamical range [58]. In our future research we will test the validity of the complete confabulation conjecture across different types of RCs, other machine learning paradigms, different tasks, and levels of biological realism.
@@ Appendix A RC design and training parameters
[…]
@@ 3. C3:
Appendix C Most common reconstruction route for $\pazocal{L}$
In Fig. 21 we show the most common way that the RC provides a reconstruction of $\pazocal{L}$ when increasing $\rho$ from 0 for one of the matrices $\textbf{M}_{i}$ from our ensemble. In each panel of Fig. 21, the black dotted curves correspond to the output of the closed-loop RC (Eq. (14)) when initialised with the last point of the training, $\bm{r}(t_{train})$. The same colour scheme from Fig. 6 is used to illustrate different UAs that we find. In panel (a) we show that while the closed-loop RC is unable to provide any reconstruction of $\pazocal{L}$, the brief transient behaviour of the RC’s state approaching UA1 resembles Lorenz-like behaviour. In panels (b)–(e) we see that before approaching UA1 or UA2, the duration of this transient activity increases and the transient also grows in its resemblance to $\pazocal{L}$. In panel (f) we no longer see this transient behaviour as the RC is able to provide a poor reconstruction of $\pazocal{L}$. However, this reconstruction has some noticeable features on the right-hand wing ($x_{1}>0$ portion of $\mathbb{P}$) that are not consistent with $\pazocal{L}$ which is why our algorithm classifies this attractor as a poor reconstruction of $\pazocal{L}$. In panels (g)–(i) we see that as $\rho$ increases, UA2 shrinks in size. Interestingly, in panel (j), after UA2 becomes unstable, this activity on the right-hand wing of the reconstructed $\pazocal{L}$ is no longer visible, and this attractor is classified as a good reconstruction of $\pazocal{L}$. The RC maintains its good reconstruction of $\pazocal{L}$ for larger values of $\rho$. Without too much speculation, one reason for this could be that since UA2 is located near $\pazocal{L}$ in this projection, it may also be in $\mathbb{S}$, and therefore the saddle which separates these states may play a role in preventing the RC from achieving a good reconstruction of $\pazocal{L}$, and after UA2 becomes unstable this enables the RC to improve its reconstruction.
In summary, Fig. 21 shows a sequence of events where we see a transition from an unstable reconstruction of $\pazocal{L}$ to a stable reconstruction of $\pazocal{L}$ with a stable UA filling in the gaps. From further investigation we find transitions from a stable UA and unstable reconstruction of $\pazocal{L}$ to an unstable UA and stable reconstruction of $\pazocal{L}$. We also find switching dynamics via transitions from a stable UA and unstable reconstruction of $\pazocal{L}$ to an unstable UA and unstable reconstruction of $\pazocal{L}$ before another transition to an unstable UA and stable reconstruction of $\pazocal{L}$. As shown in Flynn and Amann [30], when two attractors become unstable in quick succession of one another and there are no other attractors present in $\mathbb{P}$, then the state of the RC will switch between transients of both attractors in finite time and remain in either transient for different durations of time.
Appendix D Additional reconstruction routes for $\pazocal{L}$